If 14.5 mol of an idea has at 1.05 atm and 413 K, what is the volume of the gas?
1
Expert's answer
2017-10-29T11:54:06-0400
According to ideal gas law, the volume of gas is linked to the parameters given by the following equation:

pV=nRT,
where p stands for pressure, V for volume, n of the number of the moles, T for temperature and R for ideal gas constant.

Thus, it is simple to find the volume of the gas:

V=nRT/p=(14.5(mol)·0.08205746(atm L K^(-1) mol^(-1) )·413(K))/(1.05 (atm))=468 L.

Answer: The gas occupies 468 L.